# Artificial Neural Networks for Magnetoencephalography: A review of an emerging field

### v1

@@ Abstract
Magnetoencephalography (MEG) is a cutting-edge neuroimaging technique that measures the intricate brain dynamics underlying cognitive processes with an unparalleled combination of high temporal and spatial precision. MEG data analytics has always relied on advanced signal processing and mathematical and statistical tools for various tasks ranging from data cleaning to probing the signals’ rich dynamics and estimating the neural sources underlying the surface-level recordings. Like in most domains, the surge in Artificial Intelligence (AI) has led to the increased use of Machine Learning (ML) methods for MEG data classification. More recently, an emerging trend in this field is using Artificial Neural Networks (ANNs) to address many MEG-related tasks. This review provides a comprehensive overview of how ANNs are being used with MEG data from three vantage points: First, we review work that employs ANNs for MEG signal classification, i.e., for brain decoding. Second, we report on work that has used ANNs as putative models of information processing in the human brain. Finally, we examine studies that use ANNs as techniques to tackle methodological questions in MEG, including artifact correction and source estimation. Furthermore, we assess the current strengths and limitations of using ANNs with MEG and discuss future challenges and opportunities in this field. Finally, by establishing a detailed portrait of the field and providing practical recommendations for the future, this review seeks to provide a helpful reference for both seasoned MEG researchers and newcomers to the field who are interested in using ANNs to enhance the exploration of the complex dynamics of the human brain with MEG.
@@ 1 Introduction
Artificial Neural Networks (ANNs) have had an enormous impact on most research fields, especially in computer vision and natural language processing ([1]). In recent years, ANNs have become increasingly used for brain data analyses and modeling, yielding ample evidence for their added value in neuroscience research ([2, 3]). Deep Learning (DL) has been applied in the analysis of various types of brain imaging modalities, including structural Magnetic Resonance Imaging (MRI), functional Magnetic Resonance Imaging (fMRI), Electroencephalography (EEG), and Magnetoencephalography (MEG). While using ANNs specifically with EEG has been the subject of a comprehensive review (e.g. [4]), using ANNs with MEG data is less common, and the current state-of-the-art has yet to be surveyed. The present review seeks to fill in this gap by providing a detailed survey of the main applications of ANNs in MEG research to date. The strengths and limitations of the approaches reviewed allow us to highlight the potential of ANNs in future MEG research and the main challenges in this field.
@@ 1.1 Magnetoencephalography and brain imaging
MEG is a non-invasive electrophysiological brain imaging technique that detects the magnetic signals generated by the electrical activity generated by synchronized neuronal ensembles in the brain. Although it shares essential similarities with EEG (such as the electrophysiological origins of the signals and the millisecond-range temporal resolution), MEG offers unique advantages. A significant merit of MEG over EEG lies in its lesser susceptibility to spatial smearing caused by variations in the conductivity of different brain tissues and the skull. MEG is sensitive to magnetic permittivity changes, which are relatively homogeneous in the brain ([5]). As a result, MEG is a better-suited technique for source reconstruction and the analysis of spatio-temporal brain dynamics in cortical source space. Compared to fMRI, another widely used functional neuroimaging technique, MEG has several strengths. While the latter generates images of the brain only every few seconds, MEG captures brain activity on a millisecond scale. Furthermore, unlike fMRI, which measures variations in oxygenation levels in the brain, MEG provides a direct measure of the brain’s electrophysiological activity [6]. Both EEG and MEG produce sensor-level data that are highly intercorrelated and have limited spatial interpretability. These ambiguities can be reduced with MEG by applying source localization techniques ([7, 8]). Although source reconstruction does not entirely resolve the field spread problem or linear mixing issue, it does facilitate the exploration of the functional role of specific brain areas with more anatomical precision than what can be achieved with sensor-level analysis [9]. Overall, given its high temporal and spatial resolution, MEG has become an established imaging modality considered particularly suited for addressing neuroscientific inquiries involving intricate spatial, temporal, and spectral patterns of distributed brain dynamics. These include, for instance, the real-time integration of information in the brain from one or multiple modalities and the neural dynamics that underpin advanced cognitive functions such as language processing and decision-making. For a complete overview of the advantages and drawbacks of MEG in the field of neuroscience, see [10].
@@ 1.2 Standard M/EEG analysis pipeline
Generally, the M/EEG analysis pipeline consists of a sequence of processing steps that start from the sensor-level raw data and, after a series of manipulations, ultimately yield observations that address a research question. Standard preprocessing pipelines often include data cleaning and artifact rejection procedures, including channel rejection, notch-filtering, and Independent Component Analysis (ICA). M/EEG recordings yield rich and high-dimensional data, and typically, the analysis pipeline involves computing a set of variables from the raw data, also known as hand-crafted feature extraction in the case of ML analysis. Computing these variables (or features) relies on processing pipelines that can be broadly divided into two main categories: time-domain or frequency-domain analyses (which can also be jointly explored using time-frequency representations of the data). Time domain analyses typically involve the computation of time-based features such as event-related potentials (ERPs), long-range temporal correlations (LRTC) using Detrended Fluctuation Analysis (DFA), complexity measures (e.g. Hjorth parameters, Hurst exponent), fractality measures (fractal dimension) or entropy measures (approximate entropy, sample entropy, permutation entropy). Frequency domain analyses derive measures from the frequency domain using spectral analysis (e.g. Wavelet and/or Fourier transform), with typical features including the spectral power across canonical frequency bands (e.g. delta, theta, alpha, beta, gamma). In time- or frequency-domain analyses, connectivity features can be calculated using multi-channel or whole-brain source data. Typical connectivity metrics assess different associations between signals from various sources or sensors. These metrics include magnitude-squared coherence (MSCoh), Phase Locking Value (PLV), Phase-Lag Index (PLI), and weighted PLI (wPLI), to name but a few. Different features capture complementary -although sometimes partly overlapping- aspects of the neural dynamics associated with various behavioral states. Sensor-level analyses in neuroscience are informative, but source reconstruction is essential for improving the interpretation of the spatial distribution of significant effects. This step involves defining a forward model and solving the inverse problem. Effective source reconstruction allows for precise identification of neural markers in specific brain regions, enhancing our understanding of complex brain functions and informing both hypothesis validation and exploratory data-driven research.
@@ 1.3 Machine learning versus inferential statistics
A typical M/EEG neuroscience study employs either inferential statistics, using hypothesis testing methods such as t-tests or ANOVA, or statistical learning approaches, notably machine learning (ML). The choice between these approaches depends on the specific research questions, with inferential statistics often used to test predefined hypotheses about neural activity, while ML is typically applied to explore complex patterns and predictions from high-dimensional datasets. Both approaches have significantly contributed to progress in neuroscience, as detailed in [11]. Additionally, the Bayesian framework has emerged as a complementary avenue in M/EEG analysis [12], allowing hypothesis testing from a probabilistic perspective. While hypothesis testing, whether conducted through traditional methods or Bayesian approaches, has long been a staple in neuroscience, it remains the subject of ongoing debate and scrutiny (e.g. [13]). Despite its undeniable strength and utility, one significant critique of null hypothesis significance testing (NHST) is the fact that it strongly relies on a priori choices guided by the literature and expert domain knowledge and that generalization of the findings is merely assumed on the basis that the studied sample is representative of the target population. As an alternative approach to examining hand-picked variables informed by the literature, MEG researchers have begun to use a variety of data-driven approaches. These broadly include two types of approaches to relevant feature identification with hardly any a priori information: (a) Feeding a large number of hand-crafted features into a decoding framework and assessing their relative contributions to model performance, and (b) Learning the relevant features from the data themselves, a form of data-driven automatic feature extraction known as representation learning. Unlike simpler algorithms such as Support Vector Machines (SVM), Linear Discriminant Analysis (LDA), or Decision Trees and Forests, ANNs can learn and extract useful features from the data automatically through intricate computations, resulting in an optimal data representation tailored to the specific requirements of the task. Despite its conceptual appeal, representation learning comes with its limitations, in particular when it comes to interpreting the learned features. ANNs are often criticized for being ’black boxes’ compared to more classical ML algorithms, which use hand-crafted features as inputs. Enhancing the feature interpretability is an active research topic in deep representation learning.
@@ 1.4 Artificial Neural Networks in a nutshell
Understanding Artificial Neural Networks (ANNs) begins with grasping the concept of linear regression, where the output $y$ is given by $y=wx+b$, with $w$ representing the weight, $x$ the input, and $b$ the bias. In ANNs, this equation gets extended by an activation function $f$, transforming it to $y=f(wx+b)$. As the neural network gains depth through additional layers, these equations combine to form increasingly complex mathematical models. A neural network’s complexity, or capacity, indicates its ability to model intricate functions, but this can be a double-edged sword. High capacity often leads to overfitting, where the model learns the training data too well, capturing noise and inaccuracies, thus performing poorly on new data. To mitigate overfitting, regularization techniques such as Maxpooling, which reduces feature map dimensions, Dropout, which nullifies a subset of neurons during training, and Batch Normalization, which standardizes layer outputs, can be applied. Another regularization method comes in the form of Data Augmentation, where the initial dataset can be enlarged by introducing noise to the data, rotating, or cropping it. Learning in ANNs occurs through optimization, typically employing Stochastic Gradient Descent (SGD). During each training iteration, the model’s parameters are adjusted based on computed gradients from a loss function, facilitated by a technique known as backpropagation. As we progress to more advanced ANNs, we encounter Convolutional Neural Networks optimized for image data, Recurrent Neural Networks ideal for sequences, and Attention Models that allocate varying degrees of focus on different parts of the input. Overall, the domain of Artificial Neural Networks (ANNs) extends from fundamental mathematical principles to intricate architectures and sophisticated methods for effective training.
@@ 1.6 Goal of this review
This review provides a comprehensive survey of ANN applications in magnetoencephalography (MEG) data analysis, highlighting their benefits and potential added value. These techniques have proven to be remarkably effective in various neuroscience contexts. By providing examples of functional architectures and techniques, we aim to provide an up-to-date overview of the concrete progress in ANN-based MEG studies. We hope to expedite researchers’ path to developing their own implementations. More globally, we hope this review will encourage researchers well-versed with MEG to integrate deep learning techniques into their work, and conversely encourage experts in deep learning with an interest in neuroscience to consider MEG as a particularly promising brain measurement modality in the context of Neuro-AI research applications.
@@ 2.1 Litterature research
For this review, we collected English-language conference and journal papers using PubMed, Google Scholar, arXiv, and bioRxiv. We used the following query to find the papers referenced in this review: ("deep learning" OR "artificial neural network*" OR "Convolutional neural net*" OR CNN OR "Recurrent neural net*" OR RNN) AND ("MEG" OR Magnetoencephalogra*) Relevant papers were selected by examining the title first, followed by the abstract. Finally, we used a PDF search for the terms from the query to better understand how they are used. This study includes conference papers, journal papers, and pre-prints. The final inclusion criterion requires papers to possess a valid Digital Object Identifier (DOI) and a publication date preceding November 2024.
Tables 3, 4, 5, 6 provide an overview of the key information extracted from the included studies for this survey. More detailed information on each paper was collected in a table we have made available online as a Google sheet: https://tinyurl.com/ub3s5mr. See Table 1 for details on each information category extracted from the papers.
2.2 Types of studies
In conducting this review, we opted to categorize the included studies into three distinct groups: (1) Classification, (2) Modeling, and (3) Other. The ’Classification’ category was split into four sub-groups: Decoding, Brain-Computer Interfaces (BCI), Clinical Applications or Computer-Aided Diagnosis (CAD), and Event Detection. Studies under ’Decoding’ primarily leverage classification as a statistical tool to identify the involvement of specific brain features in certain tasks. The BCI subcategory includes studies that incorporate ANNs in their BCI framework. Meanwhile, the ’Clinical’ subcategory includes studies that utilize ANNs to enhance clinical diagnosis or prognosis, and the ’Event Detection’ subcategory regroups studies using ANNs to continuously detect events in MEG data. Studies in the ’Modeling’ category primarily focus on comparing activations across ANNs and the brain. The ’Other’ category gathers the studies that do not fit the aforementioned groups and includes research utilizing ANNs for preprocessing tasks such as artifact detection and removal or source localization (See Table 2 for further details).

### v3

@@ Abstract
Objective: Magnetoencephalography (MEG) is a cutting-edge neuroimaging technique that measures the intricate brain dynamics underlying cognitive processes with an unparalleled combination of high temporal and spatial precision. While MEG data analytics have traditionally relied on advanced signal processing and mathematical and statistical tools, the recent surge in artificial intelligence (AI) has led to the growing use of machine learning (ML) methods for MEG data classification. An emerging trend in this field is the use of artificial neural networks (ANNs) to address various MEG-related tasks. This review aims to provide a comprehensive overview of the state of the art in this area. Approach: This topical review included studies that applied ANNs to MEG data. Studies were sourced from PubMed, Google Scholar, arXiv, and bioRxiv using targeted search queries. The included studies were categorized into three groups: ’Classification’, ’Modeling’, and ’Other’. Key findings and trends were summarized to provide a comprehensive assessment of the field.Main results: The review identified 119 relevant studies, with 70 focused on ’Classification’, 16 on ’Modeling’, and 33 in the ’Other’ category. ’Classification’ studies addressed tasks such as brain decoding, clinical diagnostics, and BCI implementations, often achieving high predictive accuracy. ’Modeling’ studies explored the alignment between ANN activations and brain processes, offering insights into the neural representations captured by these networks. The ’Other’ category demonstrated innovative uses of ANNs for artifact correction, preprocessing, and neural source localization. Significance: By establishing a detailed portrait of the current state of the field, this review highlights the strengths and current limitations of ANNs in MEG research. It also provides practical recommendations for future work, offering a helpful reference for seasoned researchers and newcomers interested in using ANNs to explore the complex dynamics of the human brain with MEG.
@@ 1 Introduction
Artificial neural networks (ANNs) have had an enormous impact on most research fields, especially in computer vision and natural language processing ([1]). In recent years, ANNs have become increasingly used for brain data analyses and modeling, yielding ample evidence for their added value in neuroscience research ([2, 3]). Deep learning (DL) has been applied in the analysis of various types of brain imaging modalities, including structural magnetic iesonance imaging (MRI), functional magnetic resonance imaging (fMRI), electroencephalography (EEG), and magnetoencephalography (MEG). While using ANNs specifically with EEG has been the subject of comprehensive reviews (e.g. [4, 5]), using ANNs with MEG data is less common, and the current state-of-the-art has yet to be surveyed. This combination, however, is of growing interest. MEG provides direct, time-resolved recordings of brain activity, which can be used to compare the dynamics and internal representations of ANN models to those observed in the brain. For ANN researchers, this opens up possibilities for testing model outputs against high-temporal-resolution biological signals and for training models that simulate aspects of cognitive processing. For MEG researchers, artificial neural networks (ANNs) offer powerful tools to classify different brain states, uncover hidden patterns in brain activity, and model the complex structure of MEG datasets. These shared opportunities have led to an increasing number of studies exploring how MEG and ANNs can be integrated in practice. The present review seeks to fill in this gap by providing a detailed survey of the main applications of ANNs in MEG research to date. The strengths and limitations of the approaches reviewed allow us to highlight the potential of ANNs in future MEG research and the main challenges in this field.
@@ 1.1 Magnetoencephalography and brain imaging
MEG is a non-invasive electrophysiological brain imaging technique that detects magnetic signals generated by synchronized neuronal ensembles in the brain, typically recorded with approximately 100–300 sensors depending on the system. Although it shares similarities with EEG, such as electrophysiological signal origins and millisecond-range temporal resolution, MEG offers distinct advantages over EEG and other modalities like fMRI. Unlike EEG, MEG is less susceptible to spatial smearing from variations in conductivity across brain tissues and the skull, due to its sensitivity to magnetic permittivity changes, which are relatively homogeneous in the brain ([6]). This makes MEG particularly well-suited for source reconstruction and analyzing spatio-temporal brain dynamics in cortical source space ([7, 8]). In contrast to fMRI, which measures hemodynamic responses only indirectly reflecting neural activity with a temporal resolution of several seconds, MEG directly captures electrophysiological activity on a millisecond scale ([9]).
Overall, with its high temporal resolution and improved spatial precision over EEG, MEG excels in studying complex brain dynamics, despite challenges like field spread that source reconstruction mitigates but does not fully resolve ([10]). Both EEG and MEG produce sensor-level data that are highly intercorrelated and have limited spatial interpretability. These ambiguities can be reduced with MEG by applying source localization techniques ([8, 10]). Compared to fMRI, another widely used functional neuroimaging technique, MEG’s millisecond-scale resolution provides a significant advantage for capturing rapid neural events, whereas fMRI’s lower sampling rate limits its temporal precision ([9]). Although source reconstruction does not entirely resolve the field spread problem or linear mixing issue, it facilitates exploration of the functional role of specific brain areas with more anatomical precision than sensor-level analysis ([11]).
Given its high temporal and spatial resolution, MEG has become an established imaging modality particularly suited for addressing neuroscientific inquiries involving intricate spatial, temporal, and spectral patterns of distributed brain dynamics. These include the real-time integration of information from one or multiple modalities and the neural dynamics underpinning advanced cognitive functions such as language processing and decision-making. For a complete overview of the advantages and drawbacks of MEG in neuroscience, see [12].
@@ 1.2 Standard M/EEG analysis pipeline
Generally, the M/EEG analysis pipeline consists of a sequence of processing steps that start from the sensor-level raw data and, after a series of manipulations, ultimately yield observations that address a research question. Standard preprocessing pipelines often include data cleaning and artifact rejection procedures, including channel rejection, notch-filtering, and independent component analysis (ICA [13]).
M/EEG recordings yield rich and high-dimensional data, and typically, the analysis pipeline involves computing a set of variables from the raw data, also known as hand-crafted feature extraction in the case of ML analysis. Computing these variables (or features) relies on processing pipelines that can be broadly divided into two main categories: time-domain or frequency-domain analyses (which can also be jointly explored using time-frequency representations of the data ([14, 15]).
Time domain analyses typically involve the computation of features such as event-related potentials (ERPs [16]), long-range temporal correlations (LRTC) using detrended fluctuation analysis (DFA [17] ), complexity measures (e.g. Hjorth parameters [18], Hurst exponent [19], fractality measures (fractal dimension) or entropy measures (approximate entropy [20], sample entropy [21], permutation entropy [22]). Frequency-domain analyses typically employ spectral analysis methods, such as the wavelet ([23]) or Fourier transform ([24, 14]), to derive features like the spectral power across canonical frequency bands (e.g. delta, theta, alpha, beta, gamma).
In time- or frequency-domain analyses, connectivity features can be calculated using multi-channel or whole-brain source data. Typical connectivity metrics assess different associations between signals from various sources or sensors. These metrics include magnitude-squared coherence (MSCoh [25]), phase locking value (PLV [26]), phase-lag index (PLI [27]), and weighted PLI (wPLI [28]), to name but a few. Different features capture complementary -although sometimes partly overlapping- aspects of the neural dynamics associated with various behavioral states.
Sensor-level analyses in neuroscience are informative, but source reconstruction is essential for improving the interpretation of the spatial distribution of significant effects. This step involves defining a forward model and solving the inverse problem ( [29] ). Effective source reconstruction allows for precise identification of neural markers in specific brain regions, enhancing our understanding of complex brain functions and informing both hypothesis validation and exploratory data-driven research.
@@ 1.3 Machine learning versus inferential statistics
A typical M/EEG neuroscience study employs either inferential statistics, using hypothesis testing methods such as t-tests or ANOVA, or statistical learning approaches, notably machine learning (ML). The choice between these approaches depends on the specific research questions, with inferential statistics often used to test predefined hypotheses about neural activity, while ML is typically applied to explore complex patterns and predictions from high-dimensional datasets. Both approaches have significantly contributed to progress in neuroscience, as detailed in [30]. Additionally, the bayesian framework has emerged as a complementary avenue in M/EEG analysis ([31]), allowing hypothesis testing from a probabilistic perspective. While hypothesis testing, whether conducted through traditional methods or bayesian approaches, has long been a staple in neuroscience, it remains the subject of ongoing debate and scrutiny ([32]). Despite its undeniable strength and utility, null hypothesis significance testing (NHST) has several limitations. First, it heavily relies on a priori choices guided by literature and expert domain knowledge. Additionally, it assumes generalization of findings based solely on the representativeness of the studied sample to the target population. As an alternative approach to examining hand-picked variables informed by the literature, MEG researchers have begun to use a variety of data-driven approaches. These broadly include two types of approaches to relevant feature identification with hardly any a priori information: (a) Feeding a large number of hand-crafted features into a decoding framework and assessing their relative contributions to model performance, and (b) Learning the relevant features from the data themselves, a form of data-driven automatic feature extraction known as representation learning. Unlike simpler algorithms such as support vector machines (SVM), linear discriminant analysis (LDA), or decision trees which generally rely on hand-crafted features, ANNs can automatically learn and extract relevant features from data. Despite its conceptual appeal, representation learning comes with its limitations, in particular when it comes to interpreting the learned features. ANNs are often criticized for being ’black boxes’ compared to more classical ML algorithms, which use hand-crafted features as inputs. Enhancing the feature interpretability is an active research topic in deep representation learning.
@@ 1.4 Artificial neural networks in a nutshell
Understanding ANNs begins with grasping the concept of linear regression, where the output $y$ is given by $y=wx+b$, with $w$ representing the weight, $x$ the input, and $b$ the bias. In ANNs, this equation gets extended by an activation function $f$, transforming it to $y=f(wx+b)$ ([33, 34]). As the neural network gains depth through additional layers, these equations combine to form increasingly complex mathematical models. A neural network’s complexity, or capacity ([35]), indicates its ability to model intricate functions, but this can be a double-edged sword. High capacity often leads to overfitting ([35]), where the model learns the training data too well, capturing noise and inaccuracies, thus performing poorly on new data. To mitigate overfitting, regularization techniques such as maxpooling ([36]), which reduces feature map dimensions, dropout ([37]), which nullifies a subset of neurons during training, and batch normalization ([38]), which standardizes layer outputs, can be applied. Another regularization method comes in the form of data augmentation ([39]), where the initial dataset can be enlarged by introducing noise to the data, rotating, or cropping it. Learning in ANNs occurs through optimization, typically employing stochastic gradient descent (SGD [40]). During each training iteration, the model’s parameters are adjusted based on computed gradients from a loss function, facilitated by a technique known as backpropagation ([41]). As we progress to more advanced ANNs, we encounter convolutional neural networks ([42]) optimized for image data, recurrent neural networks ([43]) ideal for sequences, and attention models ([44]) that allocate varying degrees of focus on different parts of the input. Another important class includes autoencoders, which are typically unsupervised networks designed to learn efficient data codings. They consist of an encoder compressing the input to a latent representation and a decoder reconstructing the input from this representation, aiming to capture essential data features. Autoencoders can utilize various architectures like CNNs or RNNs. Sparse Autoencoders (SAEs) are a variant that adds sparsity constraints to encourage compact representations. Overall, the domain of ANNs extends from fundamental mathematical principles to intricate architectures and sophisticated methods for effective training.
While the ANNs described in this review originate from early inspiration in neuroscience, it is important to distinguish them from biologically detailed models used in computational neuroscience. Models such as integrate-and-fire neurons, Hebbian learning mechanisms, spiking neural networks, and neural mass models aim to capture the physiological behavior of neural systems with varying levels of abstraction. In contrast, the ANNs considered in this review are primarily optimized for predictive performance and scalability in ML applications. This distinction highlights the divergence between biologically motivated modeling and task-driven computational approaches — both valuable, but grounded in different research priorities ([45]).
@@ 1.6 Goal of this review
A growing number of AI researchers are turning to MEG as a rich source of non-invasive brain-wide neural data, recognizing its high spatio-temporal resolution and ability to capture complex cognitive dynamics as key assets for advancing more efficient, robust, and brain-inspired neural network models. This review provides a comprehensive survey of ANN applications in MEG data analysis, highlighting their benefits and potential added value. These techniques have proven to be shown to be useful in various neuroscience contexts. By providing examples of architectures that emulate aspects of visual processing and techniques, we aim to provide an up-to-date overview of the demonstrated the potential of this approach in ANN-based MEG studies. We hope to expedite researchers’ path to developing their own implementations. More globally, we hope this review will encourage researchers well-versed with MEG to integrate deep learning techniques into their work, and conversely encourage experts in deep learning with an interest in neuroscience to consider MEG as a particularly promising brain measurement modality in the context of Neuro-AI research applications.
@@ 2.1 Litterature research
For this review, we collected English-language conference and journal papers using PubMed, Google Scholar, arXiv, and bioRxiv. We used the following query to find the papers referenced in this review: ("deep learning" OR "artificial neural network*" OR "Convolutional neural net*" OR CNN OR "Recurrent neural net*" OR RNN) AND ("MEG" OR Magnetoencephalogra*). Relevant papers were selected by examining the title first, followed by the abstract. Finally, we used a PDF search for the terms from the query to better understand how they are used. This study includes conference papers, journal papers, and pre-prints. The final inclusion criterion requires papers to be primary research articles (i.e., not reviews), possess a valid digital object identifier (DOI) and a publication date preceding November 2024. Figure 1 provides a schematic overview of the literature search and selection process employed in this review.
Figure 1: Schematic overview of the literature review process. An initial search query across specified databases yielded over 10,000 results. These were screened based on language, DOI availability, publication date (< Nov 2024), and exclusion of reviews, resulting in 119 included studies. Following data extraction, studies were categorized into three main groups: ’Classification’ (N=70; further subdivided into decoding, BCI, clinical, and event detection), ’Modeling’ (N=16; subdivided by cortical focus), and ’Other’ methodological applications (N=33; subdivided into method, source localization, and preprocessing).
Tables 2, 3, 4, 5 provide an overview of the key information extracted from the included studies for this survey. More detailed information on each paper was collected in a table we have made available online as a Google sheet: https://tinyurl.com/ub3s5mr. See table 1 for details on each information category extracted from the papers.